OpenStudy (anonymous):

The concentration of cholestrol (C27H46O) in a patients bloodd was found to be a 6.0mmol/dm^3 What is the concentration of cholestrol in g/dm^3

OpenStudy (anonymous):

\[g C_{27}H_{46}O = 27(12.011)+46(1.00794)+(15.9994)\frac{g}{mol}=386.66164\frac{g}{mol}\] \[6.0\frac{mmol}{dm^{3}}\frac{1 mol}{1000 mmol}\frac{386.66164 g}{1 mol}=2.3\frac{g}{dm^{3}}\]

OpenStudy (anonymous):

thank u for the anwer can u plis explain plis

OpenStudy (anonymous):

There is 1 mol for every 1000 mmol. The molar mass of (C27H46O) is 386.66 per 1 mol. If you look, the mmol and mol cancel leaving dm^3 and g.